When the ferry engine is set to optimal operation mode, its speed in relation to still water is 8 km / h. What is the speed of the ferry in relation to the shore if its engine is set to the optimal mode of operation, and the ferry moves vertical to the course of the river whose speed in relation to the shore is 6 km / h?

The solution is 10 km / h (2.78 m / s) but I don't know the procedure.



1
Expert's answer
2021-03-14T19:13:04-0400
"v=\sqrt{v^2_x+v_y^2}\\v=\sqrt{8^2+6^2}=10\frac{km}{h}\\v=\frac{10}{3.6}=2.78\frac{m}{s}"